Suppose you want to buy a house that cost $100, 000.00 The bank requires a 5.0% down payment and will charge 10.80% interest compounded monthly. Suppose you put down the least amount that you could and borrow the rest from the bank. If you made monthly payments to the bank for 15 years, how much would each payment be?
Answer #1

total value of house = 1,00,00,000

down payment= 5,00,000

loan amount = 95,00,000

pv= A(1-(1+R)^-N / R

9500000=a(1-(1+10.8/12)^-12*15 / 10.8/12

9500000= A(1.009)^-180/.009

9500000= A(0.8)/.009

9500000=A(88.96)

A= 9500000/88.96

A=106789